It a 150 m alcoho ess
Answer #1

(a) V = V0(1 + \alpha*∆T) where \alpha is the volume temperature extension coefficient

so,

V = 150 ( 1 +  0.00109 * (21 - 4.8))

V = 152.6487 mL

Therefore, overflow will be

V = 2.6487 mL

---------------------------------------------------------------------------------------------------

(b) In case of water

V = 150 ( 1 +   0.000214 * 16.2)

V = 150.52002 mL

V = 0.52002 mL

less overflow of water = 2.6487 - 0.52002 = 2.12868 mL

Note that the values of expansion cofficients are taken from text book and may have slight difference from book to book. However, the method is fine.